A Complete Bibliography of Publications in Fish Physiology and Biochemistry: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

21 August 2024 Version 1.05

Title word cross-reference

(n − 3) [113]. (n − 6) [113]. ^{+} [92, 100, 140, 89]. ^{125} [120]. ^{51} [120]. ^{6} [14]. ^{9} [14]. _{3} [142, 197, 205]. α [12, 27, 119, 185, 14, 161]. β [142, 127, 146, 27, 119, 185, 14, 155, 161]. ′ [12, 78].

-ATPase [100, 140]. -ATPases [92]. -dihydroprogesterone [27]. -dihydroxy-4-pregnen-3-one [119, 185, 14]. -estradiol [127]. -hydroxy [27]. -linked [89]. -malic [114]. -methyltestosterone [12]. -triiodo-l-thyronine [78].

/K [100].

17- [12].

21-trihydroxy-4-pregnen-3-one [161].

4-dinitrophenol [18]. 450 [9, 136].

5-triiodo-L-thyronine [12].

7-alkoxycoumarins [136].

Aa [31]. ability [76]. Absorption [56]. acclimated [83]. Accumulation [17]. acid [153, 129, 115, 112, 113, 85, 194]. acid-base [85]. acids [57, 61, 68, 113]. Acipenser [99]. action [201, 27, 185, 156]. actions [196]. activation [91]. Active [11, 128]. activities [155, 176]. activity [4, 166, 175, 170, 71, 119, 52, 140, 194]. adaptation [200, 97]. added [115]. adenosine [18]. adenosine-5-triphosphate [18]. adjustments [39]. adrenaline [50, 130]. adrenergic [131]. adult [125, 177]. affect [78]. African [137, 98, 122, 124, 186, 123]. after [198, 13, 112, 123]. aglomerular [202]. air [100]. air-breathing [100]. Ala [14]. Albula [68, 84, 71]. albumin [120]. alkoxycoumarins [136]. alosoides [14]. Alpha [85]. aluminium [115]. always [6]. American [205, 10]. americanus [149, 173]. amino [61, 68]. ammocoetes [72]. Ammonia [53, 104]. among [70]. AMP [27, 109]. Anabas [155]. anabolic [129]. analogs [174]. analysis [73, 74, 97, 157, 147]. Androgen [54]. androgens [146]. angiotensin [51]. angiotensin-converting [51]. Anguilla [39, 12, 181, 10]. animal [44, 176]. animals [128]. Announcement [64, 40, 20]. Annual [80, 45, 28]. Antarctic [133]. Antifreeze [126]. antifreezes [75]. application [23]. aquaculture [23]. aquatic [121]. arginine [70, 37]. articles [87, 134]. ascorbate [36]. ascorbate-2-sulfate [36]. ascorbic [153]. aspects [93, 204]. associated [173, 62]. Atlantic [141, 140, 107, 138, 194, 101]. ATPase [100, 140]. ATPases [92]. aurata [169]. auratus [178]. Australian [206]. australis [72]. Autocatalytic [74]. Autonomic [138]. avoid [121].

baeri [99]. balance [197, 43]. base [85]. based [49]. bass [198, 80, 26]. batrachus [62]. before [123]. beta [85]. between [91, 180, 197, 102, 153, 124]. Bilirubin [96]. Bilirubin-binding [96]. binding [96]. bioactivity [206, 169]. bioassay [176]. biochemical [204, 90, 37]. Biology [32]. Bloch [108, 67, 155]. blood [149, 39, 120, 50, 85, 147]. blood-perfused [147]. blue [66, 67]. body [7, 12, 120, 56]. bonefish [68, 84]. boreal [75]. bovine [120, 13]. brain [173, 103, 153, 144, 129, 186]. brains [113]. branchial [73]. breakdown [146]. breathers [121]. breathing [100]. brook [63, 119, 160]. brown [95, 54]. buffering [42]. burn [74].

c [109]. c-AMP [109]. C. [71]. cadmium [139]. calcium [203]. cAMP [142]. canola [135, 49, 110]. capacity [16, 42]. Carassius [178]. carbohydrate [7, 132, 3]. Cardiac [86, 66, 149, 73, 133]. Cardiovascular [39]. care [46]. carnio [46]. carotenoid [96]. carotenoid-carrying [96]. carp [31, 129, 131, 83, 28, 184]. carpio [31, 129]. carps [146]. carrying [96]. catalyzed [60]. Catechol [186]. Catechol-O-methyltransferase [186]. Catecholamine [173, 103]. catecholamines [195]. Catecholestrogens [137, 186]. catfish [137, 175, 62, 98, 156, 122, 124, 186, 123]. Catla [146]. cell [188, 165, 74, 112]. cells [9, 154, 50, 187, 131]. Central [24]. cerebral [170]. chain [151]. Changes [143, 62, 200, 11, 29, 63, 46, 187, 67, 13]. Channa [108]. characteristics [157]. Characterization [92, 170, 98, 51, 84]. chicken [94]. Chinese [145]. Chinook [191, 143, 106]. chloride [115]. Chondrostei [99]. chorions [157]. chromatography [123]. chronic [132, 192]. chum [4, 96, 76, 200, 47]. Circulatory [195]. cirrhatus [150]. Cirrhinus [146]. Clarias [137, 62, 98, 122, 124, 186, 123]. classical [193]. cod [52, 43, 67, 113, 138, 130, 151]. Coho [94, 143, 44, 16, 48, 106, 79, 183, 118]. colias [67]. combinations [129]. common [46, 28]. compared [3]. Comparison [120]. compensation [135, 110]. composition [7, 12, 69]. compositions [112, 113]. concentrations [200]. condition [67]. conditions [108, 97]. Conference [40, 20]. constituents [93]. consumption [72]. continuous [82]. Control [81, 160, 181, 182, 199, 22, 193]. controlled [23]. controlling [202]. conversion [129]. converting [51]. copper [45]. coregoninae [71]. Coregonus [71]. corpuscle [75]. corpuscles [206, 204]. correct [121]. corticosteroids [146]. corticotropes [166]. Cortisol [10, 192, 194]. cost [73]. coupling [165]. Cr [120]. Cr-erythrocyte [120]. Creatine [117]. critical [147]. Ctenopharyngodon [184]. culture [168, 112]. cyanide [86]. cycle [80, 62, 45, 28]. cyclic [27]. Cyprinidae [90, 97]. Cyprinus [31, 46, 129]. Cytochrome [9, 136].

D [205, 14]. D-Ala [14]. dabryanus [145]. daily [109]. damage [38]. dealkylation [136]. death [74]. deficiency [153, 8]. degradation [4]. dehydrogenase [133, 89, 58]. deiodination [127]. Delayed [26]. dependent [9, 142]. depot [106]. derivatives [146]. determined [37, 104, 123]. developing [43]. development [23, 75, 116, 187, 190]. Developmental [93]. Dicentrarchus [80, 109, 26]. diet [109, 90, 3]. Dietary [7, 37, 70, 56, 132, 78, 153, 5]. diets [60, 49]. differences [131]. different [152, 129, 83, 26, 151]. Differentiation [187]. dihydroprogesterone [27]. dihydroxy [119, 185, 14]. dinitrophenol [18]. directions [158]. disease [192]. dissimilar [86]. Distribution [58, 57, 56, 170, 53, 186, 177]. Diurnal [108]. Does [78, 2]. Dopamine [145, 137, 186]. Drinking [52]. duality [167]. duct [160]. during [4, 73, 63, 44, 119, 140, 187, 200, 45, 139, 122, 28, 123]. dying [115].

ecological [20]. ecologically [86]. eel [206, 39, 12, 195, 204]. eels [181, 205, 10, 144]. Effect [15, 103, 48, 194, 11, 12, 60, 100, 127, 5, 129, 155, 105]. effectiveness [146]. Effects [31, 50, 49, 14, 109, 192, 174, 79, 85, 18, 175, 44, 34, 90, 3, 101]. efficiency [129]. eft [176]. egg [157]. eggs [143, 17, 43]. Elasmobranch [6, 66]. elevation [192]. embryos [71]. endocrine [162, 177]. Endocrinology [168, 77, 125, 164]. endocytose [154]. energetic [73]. Energetics [152, 141]. energy [197]. Enhanced [198]. enhancement [142]. Environmental [34, 90]. environments [194]. enzyme [51, 71, 140, 155, 114]. epinephrine [156, 85]. epithelium [160]. Eptatretus [150]. equilibria [50]. Erratum [88]. errors [121]. erythrocyte [120, 5]. erythrocytosis [54]. established [112]. estradiol [142, 127, 146, 155]. estradiol-17 [142, 146, 155]. ethology [20]. ethylamide [14]. European [144]. Evidence [161]. Evolutionary [159, 20]. excess [15]. exchange [107, 157]. excretion [56, 53]. experimentally [181]. Exposure [106, 100, 115, 139]. extreme [97].

factors [38]. facultative [100]. family [196, 156]. farming [22]. fast [148]. fasting [11]. fate [130]. Fatty [113, 112]. fed [9, 135, 110, 83, 151]. feed [34]. feedback [181]. feeding [129, 109, 3]. female [173, 181, 25, 179]. fibres [116]. fingerlings [198]. First [77]. fish [77, 168, 11, 180, 55, 100, 133, 126, 164, 57, 158, 203, 33, 116, 59, 109, 192, 162, 22, 53, 131, 86, 157, 89, 112, 113, 36, 58]. fish-egg [157]. fishes [20, 189, 75, 152, 41, 193, 172]. flagfish [103]. flavescens [182, 42]. floridae [103]. flounder [173, 117]. flow [149]. Fluctuations [28]. fluid [124]. follicles [16, 128, 101]. fontinalis [63, 119, 160]. food [129]. force [116]. Foreword [21]. form [172]. fourteen [123]. frame [201]. Free [68, 55, 61]. fresh [76]. freshwater [92, 62, 200, 58]. fry [17]. function [6, 66, 15, 190, 79]. functions [195]. Fundulus [128]. future [158].

Gadus [52, 43, 113, 138, 151]. gain [153]. gairdneri [81, 9, 7, 91, 93, 136, 70, 142, 15, 135, 102, 127, 29, 154, 111, 120, 199, 56, 50, 132, 78, 17, 27, 153, 49, 110, 51, 61, 8, 45, 139, 113, 3, 105]. gallbladder [81]. gametes [32]. gametogenesis [26]. gariepinus [137, 98, 122, 124, 186, 123]. gas [147, 123]. gender [103]. genes [159]. Genetic [22, 190]. Geotria [72]. germinal [146]. GH [191]. gill [100, 141]. gills [92, 51]. gilthead [169]. gland [49, 48]. glands [171]. glass [12]. glucagon [196, 156]. glucagon-family [196, 156]. glucose [80, 78]. glucuronides [30, 124]. glutathione [59]. Glycogenolytic [156]. glycoprotein [184]. glycoproteins [154]. GnRH [181, 175, 144, 172]. goby [100]. goldeye [14]. goldfish [204, 166, 18, 175, 170, 179, 178]. Gonadal [30, 173, 175]. gonadoliberin [181]. gonadotrophin [101]. Gonadotropic [111]. Gonadotropin [46, 128, 171, 181, 29, 175, 170, 179, 98, 145, 14, 174, 28, 169]. Gonadotropin-induced [46]. gonadotropin-releasing [175, 98, 174]. Gonadotropins [183, 167, 184]. gonadotrops [137]. gonads [181, 69]. gradients [104]. granulosa [101]. grass [184]. growth [7, 125, 159, 201, 12, 60, 102, 49, 129, 178, 35, 200, 37]. GTH [181, 111]. gut hagfish [150]. handling [205]. hatched [52]. hatching [71]. hCG [144]. head [147]. heat [7, 150]. heavy [151]. hemisphere [72]. Hemitripterus [149]. hepatic [127, 45, 109, 155]. Hepatocellular [97]. hepatocytes [156]. herring [114]. heteroclitus [128]. high [3]. hiodon [14]. homologous [94, 95]. Hormonal [63]. hormone [159, 201, 102, 29, 111, 143, 175, 49, 98, 178, 48, 200, 174]. hormones [180, 55, 203, 179, 35, 196, 169]. horn [66]. husbandry [128]. hydroxy [27]. hypocalcin [204]. hypophysectomy [44]. Hypothalamic [199, 178, 108]. hypoxia [39, 72, 61].

ide [97]. idell [184]. identity [96]. idus [90, 97]. II [183]. immature [10]. Immunohistochemical [99]. implications [159]. increment [7]. Indian [146]. indoleamine [103]. induced [46, 122, 118]. inducing [27, 185, 161]. Induction [139]. Influence [141, 116, 66, 132]. influences [42, 190]. influencing [178]. infusion [85]. inhibit [137]. inhibits [145]. injected [94]. innervation [148, 138]. insulin [80, 78, 109]. integrated [23]. Interacting [90]. Interaction [102, 33]. Interactions [70]. international [77]. interrenal [49, 193]. intersexual [189]. investigation [99]. Involvement [179]. iodide [15, 143]. ion [160, 140, 107, 157]. ion-exchange [157]. ionic [50]. Iron [60]. isocitrate [89]. isoenzymes [58]. isolated [9]. Isolation [84, 184]. isozymes [133].

Japanese [39]. japonica [39]. Jordanella [103]. Journal [87, 134]. juvenile [11, 94, 153, 200, 106, 183]. juveniles [17].

kept [26]. keratan [84]. keta [4, 96, 76, 200]. kidney [129, 141]. kinetics [94]. kisutch [94, 143, 48, 106, 79, 118].

l [78, 12, 102, 127]. L-thyroxine [127]. L. [31, 12, 80, 57, 46, 71, 52, 43, 109, 54, 107, 97, 82, 69, 26]. Labeo [146]. labrax [80, 109, 26]. Labriform [152]. Lactate [133, 58]. lamprey [72, 174]. lampreys [174, 177]. larvae [143, 52, 84]. larval [11]. lavaretus [71]. layers [101]. leptocephali [68]. LeSueur [10]. Leuciscus [90, 97]. levels [173, 181, 29, 143, 78, 103, 153, 108, 62, 129, 187, 45, 14, 67, 54]. Leydig [187]. LHRH [31, 14]. LHRH-Aa [31]. life [2]. lines [112]. linked [89]. lipid [132, 62, 106, 69]. lipids [5]. lipoprotein [96]. liver [9, 136, 10, 154, 129, 61, 196, 90]. loach [145]. Localization [203]. locomotion [152]. long [139, 112]. long-term [139, 112]. low [3]. lysine [70, 37].

major [146, 113]. male [173, 14, 174, 186]. malic [114]. mammalian [172, 112]. mannose [154]. mannose-terminated [154]. manuscripts [19]. marble [100]. marine [55, 126]. marmorata [100]. mass [123]. maternal [198]. maturation [31, 63, 27, 119, 185, 122, 161]. maturation-inducing [185, 161]. matured [181]. maturing [54]. meal [135, 49, 110]. meal-based [49]. mean [2]. measurements [37]. mechanical [105]. mechanism [27]. Mechanisms [185, 25, 193]. mediated [136]. meiosis [18]. meiotic [63]. melanotropes [166]. melanotus [90, 97]. melatonin [108]. membrane [5]. Metabolic [196, 3, 180, 102, 86]. metabolism [9, 11, 10, 155, 186]. metabolites [109]. metallothionein [139]. metallothionen [45]. metamorphosing [68, 84]. methods [152]. methylation [137]. methyltestosterone [12]. methyltransferase [186]. microsomes [136]. migrating [96]. migration [4]. minerals [70]. mirror [129]. MIS [27]. Mitochondrial [140, 114]. mobilization [106]. model [188, 165]. modulates [171]. monoamine [108]. Monoaminergic [173]. monodeiodination [127]. morhua [52, 43, 113, 138, 151]. morphofunctional [97]. Morphogenesis [177]. morphology [48]. Morphometrics [41]. motility [81, 91]. mrigala [146]. Mugiloididae [67]. Muscle [104, 148, 4, 129, 61, 116, 42, 114, 58, 151]. muscles [133]. myocardial [41]. Myosin [151]. myotomal [148, 105]. myotomal-skeletal [105]. Myoxocephalus [148].

N [14, 47]. N-terminal [47]. NAD [114, 89]. negative [6]. neglecta [133]. neural [113]. Neuropeptide [171, 170]. Neuropeptides [166]. neurotransmitter [153]. neutral [69]. newly [52]. niloticus [187]. non [193]. non-classical [193]. noradrenaline [130]. normal [144]. Notothenia [133]. Notothenioid [41]. novel [84]. NPY [171, 170]. nucleic [129]. Nutrition [35, 33]. nutrition-reproduction [33]. nutritional [38, 42, 97]. Nybelin [133].

O [186, 136]. O-dealkylation [136]. O. [143, 106]. Oncorhynchus [4, 96, 191, 94, 143, 76, 48, 200, 106, 79, 118]. one [119, 185, 14, 161]. ontogenetic [11]. oocyte [31, 18, 27, 119, 122]. oocytes [142, 146]. oral [78]. Oreochromis [187]. organic [57]. origin [172]. origins [195]. Osmoregulatory [76]. other [197, 146, 109]. output [149]. ovarian [16, 124, 101, 28, 123]. ovaries [46]. ovary [29, 119, 138]. ovine [102]. oviposition [124]. ovulation [31, 63, 182, 46, 122, 124, 123]. oxidation [60]. Oxidative [155]. Oxyeleotris [100]. Oxygen [72, 65].

P [114, 9, 136]. P-450 [136]. P-450-dependent [9]. Pacific [125, 23]. pancreas [162, 177]. parameters [90]. Paramisgurnus [145]. Parapercis [67]. parenchymal [154]. parr [44, 140, 82]. parr-smolt [44, 140]. parrotfish [13]. partial [84]. pathways [74]. pattern [148]. pentachlorophenate [132]. peptide [47]. peptides [178, 156]. Perca [182, 42]. perch [182, 42]. Performance [150]. perfused [130, 147]. pericardial [6, 66, 6]. period [16, 200]. periods [76]. periovulatory [16]. peroxidative [38]. perspective [164]. pH [116, 104]. pheromones [30]. phosphate [205]. phosphoglycerides [113]. photoperiod [118, 26].

Photoperiodic [25]. photothermal [108]. Physiological [204, 82, 60, 65]. Physiology [162, 49, 48, 207]. pikeperch [131]. pimozide [31]. pituitaries [128]. Pituitary [175, 171, 144, 98, 99, 184]. plaice [69]. Plasma [94, 123, 93, 55, 29, 80, 78, 61, 200, 14, 67, 109, 192]. platessa [69]. Platichthys [117]. Pleuronectes [69]. Polar [69]. polyunsaturated [113]. Positive [181]. possible [30]. potential [91, 44, 176]. Preface [1, 163]. pregnen [119, 185, 14, 161]. pregnene [146]. preovulatory [29, 179]. preparation [101, 147]. presence [119]. present [158]. pressure [66]. pressures [6]. prevention [38]. primitive [172]. Pro [14]. problem [74]. Processing [206]. production [197, 46]. profiles [123]. progesterone [146]. prolactin [199, 165, 44, 200]. prolactins [176]. prolonged [76, 61, 200, 85]. proopiocortin [47]. properties [105, 58]. propylthiouracil [79]. prostaglandins [182]. protease [4]. Protein [4, 83, 96, 206, 129, 3]. protein/high [3]. protein/low [3]. proteins [126, 61]. proteoglycan [84]. proteolytic [71]. Pseudopleuronectes [173]. pumps [73]. punctatus [108]. Purification [191]. purified [101].

radioimmunoassay [94, 95, 47]. Rafineque [14]. rainbow [81, 9, 7, 91, 171, 136, 159, 70, 142, 60, 25, 15, 135, 102, 127, 29, 154, 120, 199, 56, 50, 132, 78, 17, 27, 153, 49, 110, 51, 115, 61, 8, 45, 139, 113, 85, 37, 3, 105]. ratio [132]. ration [151]. raven [149]. reappraisal [148]. reared [76, 200]. received [19]. receptor [191, 175, 98, 119, 194]. receptors [111]. recombinant [94]. recrudescence [173]. red [50, 131, 176, 58]. Referees [87, 134]. reference [57]. regimes [26]. regulating [166]. Regulation [149, 27, 193, 24, 140, 169]. regulatory [203]. Relationship [91, 180, 197]. Relationships [153]. Relative [146]. release [171, 204, 199, 170, 202, 101]. releasing [175, 98, 174, 169]. Renal [75, 205]. renin [202]. reproduction [23, 25, 24, 33]. reproductive [125, 180, 103, 62, 42, 45, 67, 190]. requirement [8]. requirements [37]. reserves [102]. resistance [192]. Respiratory [115, 39, 85]. respirometry [121]. response [60, 108]. Responses [61, 29, 72, 192, 131, 86, 82]. responsive [128]. retinas [113]. review [190]. Reviewers [19]. rhythms [25]. Richardson [127, 132, 51, 8]. rohita [146]. Role [188, 38, 182, 27, 30]. Roots [164]. rostrata [10]. round [128]. Rutilus [57].

s [128, 59]. S-transferases [59]. salar [141, 140, 107, 82, 194, 101]. salines [65]. salinity [141]. Salmo [81, 9, 7, 91, 93, 136, 70, 142, 15, 135, 102, 127, 29, 154, 111, 120, 199, 56, 50, 132, 78, 17, 27, 153, 49, 110, 51, 141, 61, 8, 95, 45, 139, 54, 107, 113, 82, 3, 105, 194, 101]. salmon [4, 96, 191, 94, 29, 143, 76, 44, 16, 141, 140, 48, 200, 107, 106, 79, 183, 47, 118, 82, 207, 194, 101, 140]. salmonid [180, 23, 192]. salmonids [125, 201, 197]. Salvelinus [63, 119, 160]. Schneider [67]. sciences [2]. scorpius [148]. Screening [57]. sea [149, 80, 174, 26]. seabream [169]. Seasonal [42, 67, 175]. seatrout [161]. seawater [201, 200, 106]. Secretion [71, 142, 165, 145, 178]. selenium [17]. separated [101]. sera [112]. serial [2]. Sertoli [188]. serum [96, 120, 69]. Sex [189, 30, 67]. sexuality [22]. sharks [66]. shortening [116]. Signals [202]. significance [186]. signs [8]. silver [181, 144]. size [105]. sizes [151]. skeletal [114, 105]. skeleton [118]. skinned [116]. smolt [44, 140, 82]. Smoltification [93, 118, 79]. smolting [107]. sodium [132, 115]. soft [115]. solubility [65]. somatotropin [94]. Some [121, 32]. southern [72]. sp [68]. Sparus [169]. spawning [4, 96, 28, 26]. spawning-migrating [96]. special [57]. Species [131, 32, 30, 86]. spectrometry [123]. sperm [160]. spermatogenesis [188]. spermiation [174]. spleen [130]. spotted [161]. Squalus [188]. stage [12]. stages [29, 119]. Stannius [206, 204]. starry [117]. starved [9, 83]. states [85]. status [158, 103, 153]. steelhead [93]. stellatus [117]. Steroid [101, 27, 46, 179, 30, 124, 161]. Steroidogenesis [29, 122, 30, 174]. Steroidogenic [16]. Steroids [124, 189, 129, 67, 28, 123]. stimulates [106]. stimulating [48]. stimulation [13]. stimulus [165]. stimulus-secretion [165]. stores [104]. Stress [192, 115]. striped [198]. stroke [66]. structure [207]. Studies [100, 204, 207, 111, 199, 131, 169]. study [191, 18]. sturgeon [99]. submitted [87, 134]. substance [185]. substances [173]. sulfate [84, 36]. supplementation [49]. surge [179]. Survival [201, 198]. sustained [73]. swimmers [152]. swimming [73, 152, 82]. symposium [77]. synthesis [204, 185, 139, 83, 151]. system [190, 105]. systems [144].

T [142, 197]. taurine [57]. techniques [23, 22]. teleocalcin [207]. teleost [148, 32, 173, 55, 133, 204, 167, 108, 185, 109, 155]. Teleostei [90]. teleosts [197, 24, 30]. temperature [11, 116, 90, 26]. temperatures [83]. temporal [127]. TEP [44]. term [139, 112]. terminal [119, 47]. terminated [154]. testicular [187]. testis [188, 111]. testosterone [187, 14, 155]. testudineus [155]. theca [101]. them [121]. thiamin [8]. three [129]. throughout [16]. thyroid [15, 55, 143, 49, 48, 13, 79]. thyroid-stimulating [48]. Thyroidal [135, 110]. thyronine [12, 102, 78]. thyrotropin [13]. thyroxine [127, 44]. tilapia [92, 165, 187, 176]. time [201, 46, 109]. time-frame [201]. tissue [120, 41, 62]. tissues [38, 57, 106, 89, 113, 83]. toadfish [202]. Total [55]. toxicity [132]. tracers [120]. Transepithelial [107, 44, 176]. transfer [147]. transferases [59]. transformation [44, 140]. Transgenic [158]. transmembrane [91]. transport [92, 160]. treated [144]. treatment [198]. trihydroxy [161]. triiodo [12, 102, 78]. triiodo-L-thyronine [102]. triiodothyronine [198, 79]. triphosphate [18]. triploid [125]. tropical [55]. trout [81, 9, 7, 91, 93, 171, 136, 159, 70, 142, 60, 25, 15, 135, 102, 73, 127, 204, 29, 154, 111, 120, 63, 199, 56, 50, 132, 78, 17, 27, 153, 49, 110, 5, 51, 115, 119, 160, 61, 8, 95, 45, 139, 54, 131, 113, 85, 37, 3, 105, 147]. trunk [133, 151]. trutta [95, 54]. tryptophan [37]. tshawytscha [191, 143, 106]. tuna [74]. turnover [117]. two [159, 108, 30, 86, 176].

Ultrastructural [13]. ultrastructure [41]. underyearling [118]. Unequal [176]. uptake [143]. urine [126]. using [94, 152, 176]. Utilization [36, 78, 34].

Variations [45, 108, 109]. various [29]. velocity [116]. ventilatory [86]. vesicle [146]. Vitamin [205]. vitellogenin [95]. vitro [171, 142, 127, 29, 111, 199, 18, 146, 13, 169]. vivo [31, 29, 147, 169]. volume [66]. volumes [120].

Walbaum [173]. Water [43, 11, 76, 115]. waterborne [17, 139, 86]. weight [153]. white [61, 58, 151]. whitefish [131]. Whole [44, 120, 176]. whole-animal [176]. wild [82]. winter [173]. without [75, 115].

Year [128]. Year-round [128]. yellow [182, 42].

zinc [56, 45].

References

Leatherland:1986:P

[1] John F. Leatherland. Preface. Fish Physiology and Biochemistry, 1(1):1, January 1986. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02309587.

Hoar:1986:NSL

[2] William S. Hoar. A new serial in the life sciences. What does this mean? Fish Physiology and Biochemistry, 1(1):3–6, January 1986. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02309588.

Walton:1986:MEF

[3] Michael J. Walton. Metabolic effects of feeding a high protein/low carbohydrate diet as compared to a low protein/high carbohydrate diet to rainbow trout Salmo gairdneri. Fish Physiology and Biochemistry, 1(1):7–15, January 1986. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02309589.

Ando:1986:PDP

[4] Seiichi Ando, Mutsuo Hatano, and Kōichi Zama. Protein degradation and protease activity of chum salmon (Oncorhynchus keta) muscle during spawning migration. Fish Physiology and Biochemistry, 1 (1):17–26, January 1986. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02309590.

Leray:1986:EDL

[5] C. Leray, G. Nonnotte, and L. Nonnotte. The effect of dietary lipids on the trout erythrocyte membrane. Fish Physiology and Biochemistry, 1(1):27–35, January 1986. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02309591.

Abel:1986:EPF

[6] D. C. Abel, J. B. Graham, and R. Shabetai. Elasmobranch pericardial function. 1. Pericardial pressures are not always negative. Fish Physiology and Biochemistry, 1(2):75–83, May 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF02290207.

Beamish:1986:DCG

[7] F. W. H. Beamish, J. W. Hilton, and S. J. Slinger. Dietary carbohydrate and growth, body composition and heat increment in rainbow trout (Salmo gairdneri). Fish Physiology and Biochemistry, 1(2):85–91, May 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF02290208.

Morito:1986:TDS

[8] Caryl L. H. Morito, Debbie H. Conrad, and John W. Hilton. The thiamin deficiency signs and requirement of rainbow trout (Salmo gairdneri, Richardson). Fish Physiology and Biochemistry, 1(2):93–104, May 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF02290209.

Andersson:1986:CPD

[9] Tommy Andersson. Cytochrome P-450-dependent metabolism in isolated liver cells from fed and starved rainbow trout, Salmo gairdneri. Fish Physiology and Biochemistry, 1(2):105–111, May 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF02290210.

Foster:1986:CLM

[10] Glen D. Foster and T. W. Moon. Cortisol and liver metabolism of immature American eels, Anguilla rostrata (LeSueur). Fish Physiology and Biochemistry, 1(2):113–124, May 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF02290211.

Dabrowski:1986:AML

[11] Konrad R. Dabrowski. Active metabolism in larval and juvenile fish: ontogenetic changes, effect of water temperature and fasting. Fish Physiology and Biochemistry, 1(3):125–144, June 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF02290254.

Degani:1986:ETT

[12] Gad Degani and Carlos Dosoretz. The effect of 3,3′, 5-triiodo-L-thyronine and 17-α-methyltestosterone on growth and body composition of the glass stage of the eel (Anguilla anguilla L.). Fish Physiology and Biochemistry, 1(3):145–151, June 1986. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02290255.

Smith:1986:UCP

[13] Carol Johnson Smith and E. Gordon Grau. Ultrastructural changes in the parrotfish thyroid after in vitro stimulation with bovine thyrotropin. Fish Physiology and Biochemistry, 1(3):153–162, June 1986. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02290256.

Pankhurst:1986:EDA

[14] N. W. Pankhurst, G. van der Kraak, and B. Breton. Effects of (D-Ala^{6}, Pro^{9} N ethylamide) — LHRH on plasma levels of gonadotropin, 17α, 20β-dihydroxy-4-pregnen-3-one and testosterone in male goldeye (hiodon alosoides rafineque). Fish Physiology and Biochemistry, 1(3): 163–170, June 1986. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02290257.

Eales:1986:EEI

[15] J. G. Eales, Daniel G. Cyr, and Robert F. Cook. Effect of excess iodide on thyroid function of rainbow trout, Salmo gairdneri. Fish Physiology and Biochemistry, 1(4):171–177, September 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF02311133.

Kraak:1986:SCC

[16] Glen Van Der Kraak and Edward M. Donaldson. Steroidogenic capacity of Coho salmon ovarian follicles throughout the periovulatory period. Fish Physiology and Biochemistry, 1(4):179–186, September 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02311134.

Hodson:1986:AWS

[17] P. V. Hodson, J. W. Hilton, and S. J. Slinger. Accumulation of waterborne selenium by rainbow trout (Salmo gairdneri), eggs, fry and juveniles. Fish Physiology and Biochemistry, 1(4):187–196, September 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02311135.

Habibi:1986:SGO

[18] Hamid R. Habibi and Charles A. Lessman. A study of goldfish oocyte meiosis in vitro: effects of 2,4-dinitrophenol and adenosine-5-triphosphate. Fish Physiology and Biochemistry, 1(4):197–205, September 1986. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF02311136.

Anonymous:1986:RMR

[19] Anonymous. Reviewers of manuscripts received in 1985. Fish Physiology and Biochemistry, 1(4):207–208, September 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF02311137.

Anonymous:1986:CAE

[20] Anonymous. Conference announcement ecological and evolutionary ethology of fishes. Fish Physiology and Biochemistry, 1(4):209–210, September 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF02311138.

Planas:1986:F

[21] J. Planas, M. Carrillo, and S. Zanuy. Foreword. Fish Physiology and Biochemistry, 2(1-4):1, October 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF02264068.

Purdom:1986:GTC

[22] C. E. Purdom. Genetic techniques for control of sexuality in fish farming. Fish Physiology and Biochemistry, 2(1-4):3–8, October 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02264069.

Donaldson:1986:IDA

[23] Edward M. Donaldson. The integrated development and application of controlled reproduction techniques in Pacific salmonid aquaculture. Fish Physiology and Biochemistry, 2(1-4):9–24, October 1986. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02264070.

Kah:1986:CRR

[24] Olivier Kah. Central regulation of reproduction in teleosts. Fish Physiology and Biochemistry, 2(1-4):25–34, October 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/BF02264071.

Duston:1986:PMR

[25] James Duston and Niall Bromage. Photoperiodic mechanisms and rhythms of reproduction in the female rainbow trout. Fish Physiology and Biochemistry, 2(1-4):35–51, October 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF02264072.

Zanuy:1986:DGS

[26] S. Zanuy, M. Carrillo, and F. Ruiz. Delayed gametogenesis and spawning of sea bass (Dicentrarchus labrax L.) kept under different photoperiod and temperature regimes. Fish Physiology and Biochemistry, 2(1-4):53–63, October 1986. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02264073.

Jalabert:1986:ROM

[27] B. Jalabert and B. Finet. Regulation of oocyte maturation in the rainbow trout, Salmo gairdneri: role of cyclic AMP in the mechanism of action of the maturation inducing steroid (MIS), 17 α-hydroxy, 20 β-dihydroprogesterone. Fish Physiology and Biochemistry, 2(1-4): 65–74, October 1986. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02264074.

Yaron:1986:FGO

[28] Zvi Yaron and Berta Levavi-Zermonsky. Fluctuations in gonadotropin and ovarian steroids during the annual cycle and spawning of the common carp.

Fish Physiology and Biochemistry, 2(1-4):75–86, October 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02264075.

Fostier:1986:SRT

[29] Alexis Fostier and Bernard Jalabert. Steroidogenesis in rainbow trout (Salmo gairdneri) at various preovulatory stages: changes in plasma hormone levels and in vivo and in vitro responses of the ovary to salmon gonadotropin. Fish Physiology and Biochemistry, 2(1-4):87–99, October 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF02264076.

Lambert:1986:GSP

[30] J. G. D. Lambert, R. van den Hurk, and P. G. W. J. van Oordt. Gonadal steroidogenesis and the possible role of steroid glucuronides as sex pheromones in two species of teleosts. Fish Physiology and Biochemistry, 2(1-4):101–107, October 1986. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02264077.

Bieniarz:1986:EPL

[31] K. Bieniarz, P. Epler, and M. Sokołowska. Effects of pimozide and LHRHaa on carp (Cyprinus carpio L.) oocyte maturation and ovulation in vivo. Fish Physiology and Biochemistry, 2(1-4):109–114, October 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02264078.

Billard:1986:BGS

[32] R. Billard, R. Christen, and A. Saad. Biology of the gametes of some teleost species. Fish Physiology and Biochemistry, 2(1-4):115– 120, October 1986. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02264079.

Luquet:1986:INR

[33] P. Luquet and T. Watanabe. Interaction “nutrition-reproduction” in fish. Fish Physiology and Biochemistry, 2(1-4):121–129, October 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02264080.

Kaushik:1986:EEF

[34] Sadasivam J. Kaushik. Environmental effects on feed utilization. Fish Physiology and Biochemistry, 2(1-4):131–140, October 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02264081.

Matty:1986:NHG

[35] A. J. Matty. Nutrition, hormones and growth. Fish Physiology and Biochemistry, 2(1-4):141–150, October 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF02264082.

Tucker:1986:UAS

[36] Barbee W. Tucker and John E. Halver. Utilization of ascorbate-2-sulfate in fish. Fish Physiology and Biochemistry, 2(1-4):151–160, October 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF02264083.

Walton:1986:DRR

[37] M. J. Walton, C. B. Cowey, and J. W. Adron. Dietary requirements of rainbow trout for tryptophan, lysine and arginine determined by growth and biochemical measurements. Fish Physiology and Biochemistry, 2(1-4):161–169, October 1986. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02264084.

Cowey:1986:RNF

[38] C. B. Cowey. The role of nutritional factors in the prevention of peroxidative damage to tissues. Fish Physiology and Biochemistry, 2(1-4): 171–178, October 1986. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02264085.

Chan:1986:CRB

[39] D. K. O. Chan. Cardiovascular, respiratory, and blood adjustments to hypoxia in the Japanese eel, Anguilla japonica. Fish Physiology and Biochemistry, 2(1-4):179–193, October 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF02264086.

Anonymous:1986:CA

[40] Anonymous. Conference announcement. Fish Physiology and Biochemistry, 2(1-4):197–198, October 1986. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF02264087.

Johnston:1987:MUM

[41] Ian A. Johnston and Paul Harrison. Morphometrics and ultrastructure of myocardial tissue in notothenioid fishes. Fish Physiology and Biochemistry, 3(1):1–6, January 1987. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02183988.

Nelson:1987:SRN

[42] Jay A. Nelson and John J. Magnuson. Seasonal, reproductive, and nutritional influences on muscle “buffering capacity” in yellow perch (Perca flavescens). Fish Physiology and Biochemistry, 3(1):7–16, January 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF02183989.

Mangor-Jensen:1987:WBD

[43] A. Mangor-Jensen. Water balance in developing eggs of the cod Gadus morhua L. Fish Physiology and Biochemistry, 3(1):17–24, January 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF02183990.

Iwata:1987:WAT

[44] Munehico Iwata, Richard S. Nishioka, and Howard A. Bern. Whole animal transepithelial potential (TEP) of Coho salmon during the parr-smolt transformation and effects of thyroxine, prolactin and hypophysectomy. Fish Physiology and Biochemistry, 3(1):25–38, January 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02183991.

Olsson:1987:VHM

[45] Per-Erik Olsson, Carl Haux, and Lars Förlin. Variations in hepatic metallothionen, zinc and copper levels during an annual reproductive cycle in rainbow trout, Salmo gairdneri. Fish Physiology and Biochemistry, 3(1):39–47, January 1987. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02183992.

Kime:1987:GIC

[46] David E. Kime and Krzysztof Bieniarz. Gonadotropin-induced changes in steroid production by ovaries of the common care Cyprinus carnio L. around the time of ovulation. Fish Physiology and Biochemistry, 3 (1):49–52, January 1987. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02183993.

Takahashi:1987:RTP

[47] Akiyoshi Takahashi, Hiroshi Kawauchi, and Tetsuya Hirano. A radioimmunoassay for N-terminal peptide of chum salmon proopiocortin. Fish Physiology and Biochemistry, 3(2):53–61, March 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/BF02182999.

Nishioka:1987:ETS

[48] Richard S. Nishioka, E. Gordon Grau, and Howard A. Bern. Effect of thyroid-stimulating hormone on the physiology and morphology of the thyroid gland in Coho salmon, Oncorhynchus kisutch. Fish Physiology and Biochemistry, 3(2):63–71, March 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF02183000.

Leatherland:1987:ETH

[49] J. F. Leatherland, J. W. Hilton, and S. J. Slinger. Effects of thyroid hormone supplementation of canola meal-based diets on growth, and interrenal and thyroid gland physiology of rainbow trout (Salmo gairdneri). Fish Physiology and Biochemistry, 3(2):73–82, March 1987. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02183001.

Heming:1987:EAI

[50] Thomas A. Heming, David J. Randall, and Madeleine M. Mazeaud. Effects of adrenaline on ionic equilibria in red blood cells of rainbow trout (Salmo gairdneri). Fish Physiology and Biochemistry, 3(2):83–90, March 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF02183002.

Lipke:1987:CAC

[51] D. W. Lipke, R. L. Thomas, and K. R. Olson. Characterization of angiotensin-converting enzyme in the gills of rainbow trout, Salmo gairdneri (Richardson). Fish Physiology and Biochemistry, 3(2):91–97, March 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF02183003.

Mangor-Jensen:1987:DAN

[52] A. Mangor-Jensen and G. R. Adoff. Drinking activity of the newly hatched larvae of cod Gadus morhua L. Fish Physiology and Biochemistry, 3 (2):99–103, March 1987. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02183004.

Randall:1987:ADE

[53] David J. Randall and Patricia A. Wright. Ammonia distribution and excretion in fish. Fish Physiology and Biochemistry, 3(3):107–120, May 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF02180412.

Pottinger:1987:ALE

[54] T. G. Pottinger and A. D. Pickering. Androgen levels and erythrocytosis in maturing brown trout, Salmo trutta L. Fish Physiology and Biochemistry, 3(3):121–126, May 1987. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02180413.

Eales:1987:TFT

[55] J. G. Eales and S. Shostak. Total and free thyroid hormones in plasma of tropical marine teleost fish. Fish Physiology and Biochemistry, 3 (3):127–131, May 1987. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02180414.

Hardy:1987:ABD

[56] Ronald W. Hardy, Craig V. Sullivan, and Ann M. Koziol. Absorption, body distribution, and excretion of dietary zinc by rainbow trout (Salmo gairdneri). Fish Physiology and Biochemistry, 3(3):133–143, May 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF02180415.

Haller:1987:SOA

[57] T. Haller and R. Lackner. Screening for organic acids in fish tissues with special reference to the distribution of taurine in Rutilus rutilus L. Fish Physiology and Biochemistry, 3(3):145–149, May 1987. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02180416.

Wieser:1987:DPL

[58] W. Wieser, R. Lackner, and U. Platzer. Distribution and properties of lactate dehydrogenase isoenzymes in red and white muscle of freshwater fish. Fish Physiology and Biochemistry, 3(3):151–162, May 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02180417.

Nimmo:1987:GTF

[59] I. A. Nimmo. The glutathione S-transferases of fish. Fish Physiology and Biochemistry, 3(4):163–172, June 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF02180277.

Desjardins:1987:ICO

[60] L. M. Desjardins, B. D. Hicks, and J. W. Hilton. Iron catalyzed oxidation of trout diets and its effect on the growth and physiological response of rainbow trout. Fish Physiology and Biochemistry, 3(4):173–182, June 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF02180278.

Medale:1987:RPH

[61] Françoise Medale, J. P. Parent, and Françoise Vellas. Responses to prolonged hypoxia by rainbow trout (Salmo gairdneri) I. Free amino acids and proteins in plasma, liver and white muscle. Fish Physiology and Biochemistry, 3(4):183–189, June 1987. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02180279.

Lal:1987:CTL

[62] Bechan Lal and T. P. Singh. Changes in tissue lipid levels in the freshwater catfish Clarias batrachus associated with the reproductive cycle. Fish Physiology and Biochemistry, 3(4):191–201, June 1987. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02180280.

Goetz:1987:HCD

[63] Frederick W. Goetz, Alexis Y. Fostier, and Bernard Jalabert. Hormonal changes during meiotic maturation and ovulation in the brook trout (Salvelinus fontinalis). Fish Physiology and Biochemistry, 3(4): 203–211, June 1987. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02180281.

Anonymous:1987:A

[64] Anonymous. Announcement. Fish Physiology and Biochemistry, 3 (4):ii, June 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02180282.

Graham:1987:SOP

[65] Mark Graham. The solubility of oxygen in physiological salines. Fish Physiology and Biochemistry, 4(1):1–4, July 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF02073860.

Abel:1987:EPF

[66] D. C. Abel, W. R. Lowell, and R. Shabetai. Elasmobranch pericardial function 2. The influence of pericardial pressure on cardiac stroke volume in horn sharks and blue sharks. Fish Physiology and Biochemistry, 4(1):5–14, July 1987. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02073861.

Pankhurst:1987:SCR

[67] N. W. Pankhurst and A. M. Conroy. Seasonal changes in reproductive condition and plasma levels of sex steroids in the blue cod, Parapercis colias (Bloch and Schneider) (Mugiloididae). Fish Physiology and Biochemistry, 4(1):15–26, July 1987. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02073862.

Pfeiler:1987:FAA

[68] Edward Pfeiler. Free amino acids in metamorphosing bonefish (Albula sp.) leptocephali. Fish Physiology and Biochemistry, 4(1):27–36, July 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF02073863.

White:1987:PNL

[69] Ann White and Thelma C. Fletcher. Polar and neutral lipid composition of the gonads and serum of the plaice, Pleuronectes platessa L. Fish Physiology and Biochemistry, 4(1):37–43, July 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF02073864.

Chiu:1987:IAD

[70] Yvonne N. Chiu, Richard E. Austic, and Gary L. Rumsey. Interactions among dietary minerals, arginine and lysine in rainbow trout (Salmo gairdneri). Fish Physiology and Biochemistry, 4(1):45–55, July 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02073865.

Luczynski:1987:SHE

[71] M. Luczynski, J. Strzczek, and P. Brzuzan. Secretion of hatching enzyme and its proteolytic activity in coregoninae (Coregonus albula L. and C. lavaretus L.) embryos. Fish Physiology and Biochemistry, 4(2): 57–62, September 1987. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02044314.

Galloway:1987:OCR

[72] R. Galloway, I. C. Potter, and R. W. Hilliard. Oxygen consumption and responses to hypoxia of ammocoetes of the southern hemisphere lamprey Geotria australis. Fish Physiology and Biochemistry, 4 (2):63–72, September 1987. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02044315.

Farrell:1987:AEC

[73] A. P. Farrell and J. F. Steffensen. An analysis of the energetic cost of the branchial and cardiac pumps during sustained swimming in trout. Fish Physiology and Biochemistry, 4(2):73–79, September 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02044316.

Hochachka:1987:APC

[74] P. W. Hochachka and R. W. Brill. Autocatalytic pathways to cell death: A new analysis of the tuna burn problem. Fish Physiology and Biochemistry, 4(2):81–87, September 1987. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02044317.

Eastman:1987:RCD

[75] J. T. Eastman, R. B. Boyd, and A. L. DeVries. Renal corpuscle development in boreal fishes with and without antifreezes. Fish Physiology and Biochemistry, 4(2):89–100, September 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF02044318.

Hasegawa:1987:OAC

[76] S. Hasegawa, T. Hirano, and S. Arai. Osmoregulatory ability of chum salmon, Oncorhynchus keta, reared in fresh water for prolonged periods. Fish Physiology and Biochemistry, 4(2):101–110, September 1987. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF02044319.

Anonymous:1987:FIS

[77] Anonymous. First international symposium on fish endocrinology. Fish Physiology and Biochemistry, 4(2):111, September 1987. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02044320.

Hilton:1987:DOP

[78] J. W. Hilton, E. M. Plisetskaya, and J. F. Leatherland. Does oral 3,5,3 ′-triiodo-l-thyronine affect dietary glucose utilization and plasma insulin levels in rainbow trout (Salmo gairdneri)? Fish Physiology and Biochemistry, 4(3):113–120, October 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF02110878.

Sullivan:1987:ETP

[79] Craig V. Sullivan, Douglas S. Darling, and Walton W. Dickhoff. Effects of triiodothyronine and propylthiouracil on thyroid function and smoltification of Coho salmon (Oncorhynchus kisutch). Fish Physiology and Biochemistry, 4(3):121–135, October 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF02110879.

Gutierrez:1987:ACP

[80] J. Gutiérrez, J. Fernández, and J. Planas. Annual cycle of plasma insulin and glucose of sea bass. Dicentrarchus labrax, L. Fish Physiology and Biochemistry, 4(3):137–141, October 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF02110880.

Aldman:1987:CGM

[81] Göran Aldman and Susanne Holmgren. Control of gallbladder motility in the rainbow trout, Salmo gairdneri. Fish Physiology and Biochemistry, 4 (3):143–155, October 1987. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02110881.

Virtanen:1987:PRC

[82] Erkki Virtanen and Leena Forsman. Physiological responses to continuous swimming in wild salmon (Salmo salar L.) parr and smolt. Fish Physiology and Biochemistry, 4(3):157–163, October 1987. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF02110882.

Watt:1988:PST

[83] P. W. Watt, P. A. Marshall, and G. Goldspink. Protein synthesis in tissues of fed and starved carp, acclimated to different temperatures. Fish Physiology and Biochemistry, 4(4):165–173, January 1988. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF01871743.

Pfeiler:1988:IPC

[84] Edward Pfeiler. Isolation and partial characterization of a novel keratan sulfate proteoglycan from metamorphosing bonefish (Albula) larvae. Fish Physiology and Biochemistry, 4(4):175–187, January 1988. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF01871744.

Vermette:1988:EPE

[85] M. G. Vermette and S. F. Perry. Effects of prolonged epinephrine infusion on blood respiratory and acid-base states in the rainbow trout: Alpha and beta effects. Fish Physiology and Biochemistry, 4(4):189–202, January 1988. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF01871745.

Sawyer:1988:CVM

[86] Paul L. Sawyer and Alan G. Heath. Cardiac, ventilatory and metabolic responses of two ecologically dissimilar species of fish to waterborne cyanide. Fish Physiology and Biochemistry, 4(4):203–219, January 1988. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF01871746.

Anonymous:1988:RAS

[87] Anonymous. Referees of articles submitted to the journal in 1987. Fish Physiology and Biochemistry, 4(4):221–222, January 1988. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF01871747.

Anonymous:1988:E

[88] Anonymous. Erratum. Fish Physiology and Biochemistry, 4(4):223, January 1988. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01871748.

Storey:1988:NLI

[89] Kenneth B. Storey and Jeremy H. A. Fields. NAD^{+}-linked isocitrate dehydrogenase in fish tissues. Fish Physiology and Biochemistry, 5(1):1–8, March 1988. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01874723.

Rafael:1988:IED

[90] Johannes Rafael and Thomas Braunbeck. Interacting effects of diet and environmental temperature on biochemical parameters in the liver of Leuciscus idus melanotus (Cyprinidae: Teleostei). Fish Physiology and Biochemistry, 5(1):9–19, March 1988. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF01874724.

Blaber:1988:RBT

[91] Andrew P. Blaber and F. Ross Hallett. Relationship between transmembrane potential and activation of motility in rainbow trout (Salmo gairdneri). Fish Physiology and Biochemistry, 5(1):21–30, March 1988. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF01874725.

Balm:1988:CTA

[92] Paul Balm, Nico Goossen, and Sjoerd Wendelaar Bonga. Characterization of transport Na^{+}-ATPases in gills of freshwater tilapia. Fish Physiology and Biochemistry, 5(1):31–38, March 1988. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF01874726.

Bradley:1988:SST

[93] Terence M. Bradley and A. W. Rourke. Smoltification in steelhead trout (Salmo gairdneri): Developmental aspects of plasma constituents. Fish Physiology and Biochemistry, 5(1):39–47, March 1988. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/BF01874727.

Down:1988:PKI

[94] N. E. (Ted) Down, Helen M. Dye, and Lawrence M. Souza. Plasma kinetics of injected recombinant chicken somatotropin in juvenile Coho salmon (Oncorhynchus kisutch) using a homologous radioimmunoassay. Fish Physiology and Biochemistry, 5(2):49–57, May 1988. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF01875642.

Norberg:1988:HRB

[95] Birgitta Norberg and Carl Haux. An homologous radioimmunoassay for brown trout (Salmo trutta) vitellogenin. Fish Physiology and Biochemistry, 5(2):59–68, May 1988. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875643.

Ando:1988:BBP

[96] Seiichi Ando and Mutsuo Hatano. Bilirubin-binding protein in the serum of spawning-migrating chum salmon, Oncorhynchus keta: Its identity with carotenoid-carrying lipoprotein. Fish Physiology and Biochemistry, 5(2):69–78, May 1988. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875644.

Segner:1988:HAE

[97] Helmut Segner and Thomas Braunbeck. Hepatocellular adaptation to extreme nutritional conditions in ide, Leuciscus idus melanotus L. (Cyprinidae). A morphofunctional analysis. Fish Physiology and Biochemistry, 5(2):79–97, May 1988. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875645.

Leeuw:1988:CRG

[98] R. De Leeuw, P. M. Conn, and P. G. W. J. Van Oordt. Characterization of the receptor for gonadotropin-releasing hormone in the pituitary of the African catfish, Clarias gariepinus. Fish Physiology and Biochemistry, 5(2):99–107, May 1988. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875646.

Pelissero:1988:IIP

[99] C. Pelissero, J. Nunez-Rodriguez, and O. Kah. Immunohistochemical investigation of the pituitary of the sturgeon (Acipenser baeri, Chondrostei). Fish Physiology and Biochemistry, 5(3):109–119, July 1988. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF01875699.

Fenwick:1988:SEA

[100] J. C. Fenwick and T. J. Lam. Studies on the effect of air exposure on gill Na^{+} /K^{+}-ATPase of the marble goby, Oxyeleotris marmorata, a facultative air-breathing fish. Fish Physiology and Biochemistry, 5 (3):121–130, July 1988. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875700.

Wright:1988:SRS

[101] R. S. Wright and Wei xin Zhao. Steroid release from separated theca and granulosa layers of Atlantic salmon (Salmo salar) ovarian follicles: the effects of a purified salmon gonadotrophin preparation. Fish Physiology and Biochemistry, 5(3):131–139, July 1988. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF01875701.

Farbridge:1988:IBO

[102] K. J. Farbridge and J. F. Leatherland. Interaction between ovine growth hormone and triiodo-L-thyronine on metabolic reserves of rainbow trout, Salmo gairdneri. Fish Physiology and Biochemistry, 5(3):141–151, July 1988. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF01875702.

Holdway:1988:EGR

[103] D. A. Holdway, B. D. Sloley, and D. G. Dixon. Effect of gender and reproductive status on brain catecholamine and indoleamine levels of flagfish (Jordanella floridae). Fish Physiology and Biochemistry, 5(3):153–158, July 1988. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875703.

Wright:1988:MAS

[104] Patricia A. Wright and Chris M. Wood. Muscle ammonia stores are not determined by pH gradients. Fish Physiology and Biochemistry, 5 (3):159–162, July 1988. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875704.

Webb:1988:ESM

[105] P. W. Webb and C. L. Johnsrude. The effect of size on the mechanical properties of the myotomal-skeletal system of rainbow trout (Salmo gairdneri). Fish Physiology and Biochemistry, 5(3):163–171, July 1988. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF01875705.

Sheridan:1988:ESS

[106] Mark A. Sheridan. Exposure to seawater stimulates lipid mobilization from depot tissues of juvenile Coho (Oncorhynchus kisutch) and Chinook (O. tshawytscha) salmon. Fish Physiology and Biochemistry, 5 (4):173–180, October 1988. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01874793.

Primmett:1988:TIE

[107] D. R. N. Primmett, F. B. Eddy, and J. E. Thorpe. Transepithelial ion exchange in smolting Atlantic salmon (Salmo salar L.). Fish Physiology and Biochemistry, 5(4):181–186, October 1988. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF01874794.

Khan:1988:DVH

[108] I. A. Khan and K. P. Joy. Diurnal variations in hypothalamic monoamine levels in the teleost Channa punctatus (Bloch) in response to melatonin under two photothermal conditions. Fish Physiology and Biochemistry, 5 (4):187–190, October 1988. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01874795.

Perez:1988:EDF

[109] J. Perez, S. Zanuy, and M. Carrillo. Effects of diet and feeding time on daily variations in plasma insulin, hepatic c-AMP and other metabolites in a teleost fish, Dicentrarchus labrax L. Fish Physiology and Biochemistry, 5(4):191–197, October 1988. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01874796.

Leatherland:1988:TCR

[110] J. F. Leatherland and J. W. Hilton. Thyroidal compensation in rainbow trout (Salmo gairdneri) fed canola meal. Fish Physiology and Biochemistry, 5(4):199–207, October 1988. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF01874797.

Gac:1988:GHG

[111] Florence Le Gac, Bernard Breton, and Muriel Bougoussa. Gonadotropic hormone (GtH) receptors in the testis of the trout Salmo gairdneri: in vitro studies. Fish Physiology and Biochemistry, 5(4):209–217, October 1988. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF01874798.

Tocher:1988:FACa

[112] D. R. Tocher, J. R. Sargen, and G. N. Frerichs. The fatty acid compositions of established fish cell lines after long-term culture in mammalian sera. Fish Physiology and Biochemistry, 5(4):219–227, October 1988. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF01874799.

Tocher:1988:FACb

[113] Douglas R. Tocher and Douglas G. Harvie. Fatty acid compositions of the major phosphoglycerides from fish neural tissues; (n − 3) and (n − 6) polyunsaturated fatty acids in rainbow trout (Salmo gairdneri) and cod (Gadus morhua) brains and retinas. Fish Physiology and Biochemistry, 5(4):229–239, October 1988. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01874800.

Skorkowski:1988:MNP

[114] Edward F. Skorkowski and Kenneth B. Storey. Mitochondrial NAD(P)- malic enzyme from herring skeletal muscle. Fish Physiology and Biochemistry, 5(4):241–248, October 1988. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF01874801.

Malte:1988:RSR

[115] Hans Malte and Roy E. Weber. Respiratory stress in rainbow trout dying from aluminium exposure in soft, acid water, with or without added sodium chloride. Fish Physiology and Biochemistry, 5(4):249–256, October 1988. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF01874802.

Mutungi:1988:IPT

[116] Gabriel Mutungi and Ian A. Johnston. Influence of pH and temperature on force development and shortening velocity in skinned muscle fibres from fish. Fish Physiology and Biochemistry, 5(4):257–262, October 1988. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF01874803.

Danulat:1989:CTS

[117] Eva Danulat and Peter W. Hochachka. Creatine turnover in the starry flounder, Platichthys stellatus. Fish Physiology and Biochemistry, 6 (1):1–9, January 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875599.

Thorarensen:1989:SIL

[118] Helgi Thorarensen and W. Craig Clarke. Smoltification induced by a ‘skeleton’ photoperiod in underyearling Coho salmon (Oncorhynchus kisutch). Fish Physiology and Biochemistry, 6(1):11–18, January 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF01875600.

Maneckjee:1989:PDP

[119] A. Maneckjee, M. Weisbart, and D. R. Idler. The presence of 17α, 20βdihydroxy-4-pregnen-3-one receptor activity in the ovary of the brook trout, Salvelinus fontinalis, during terminal stages of oocyte maturation. Fish Physiology and Biochemistry, 6(1):19–38, January 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF01875601.

Gingerich:1989:CWB

[120] W. H. Gingerich and R. A. Pityer. Comparison of whole body and tissue blood volumes in rainbow trout (Salmo gairdneri) with ^{125}I bovine serum albumin and ^{51}Cr-erythrocyte tracers. Fish Physiology and Biochemistry, 6(1):39–47, January 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875602.

Steffensen:1989:SER

[121] John Fleng Steffensen. Some errors in respirometry of aquatic breathers: How to avoid and correct for them. Fish Physiology and Biochemistry, 6(1):49–59, January 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02995809.

Schoonen:1989:SDI

[122] W. G. E. J. Schoonen, J. G. D. Lambert, and P. G. W. J. Van Oordt. Steroidogenesis during induced oocyte maturation and ovulation in the African catfish, Clarias gariepinus. Fish Physiology and Biochemistry, 6(2):61–78, March 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875480.

vanDam:1989:PPF

[123] G. H. van Dam, W. G. E. J. Schoonen, and P. G. W. J. van Oordt. Plasma profiles of fourteen ovarian steroids before, during and after ovulation in African catfish, Clarias gariepinus, determined by gas chromatography and mass spectrometry. Fish Physiology and Biochemistry, 6(2):79–89, March 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875481.

Schoonen:1989:SSG

[124] W. G. E. J. Schoonen, J. C. M. Granneman, and J. G. D. Lambert. Steroids and steroid glucuronides in the ovarian fluid of the African catfish, Clarias gariepinus, between ovulation and oviposition. Fish Physiology and Biochemistry, 6(2):91–112, March 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF01875482.

Benfey:1989:GRE

[125] Tillmann J. Benfey, Helen M. Dye, and Edward M. Donaldson. The growth and reproductive endocrinology of adult triploid Pacific salmonids. Fish Physiology and Biochemistry, 6(2):113–120, March 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF01875483.

Fletcher:1989:APU

[126] Garth L. Fletcher, Madonna J. King, and Margaret A. Shears. Antifreeze proteins in the urine of marine fish. Fish Physiology and Biochemistry, 6(2):121–127, March 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875484.

Flett:1989:EVH

[127] P. A. Flett and J. F. Leatherland. In vitro hepatic monodeiodination of Lthyroxine and the temporal effect of 17β-estradiol on deiodination in rainbow trout (Salmo gairdneri Richardson). Fish Physiology and Biochemistry, 6(3):129–138, May 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01874769.

Lin:1989:EFH

[128] Yu-Wai Peter Lin, Mark S. Greeley, Jr., and Robin A. Wallace. Fundulus heteroclitus gonadotropin(s) 2. Year-round husbandry of animals with active pituitaries and responsive follicles. Fish Physiology and Biochemistry, 6(3):139–148, May 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01874770.

Lone:1989:EFT

[129] Khalid P. Lone. The effect of feeding three anabolic steroids in different combinations on the growth, food conversion efficiency and protein and nucleic acid levels of liver, kidney, brain and muscle of mirror carp (Cyprinus carpio). Fish Physiology and Biochemistry, 6(3):149–156, May 1989.

CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF01874771.

Ungell:1989:FNA

[130] Anna-Lena Ungell. The fate of noradrenaline and adrenaline in the perfused cod spleen. Fish Physiology and Biochemistry, 6(3):157–165, May 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF01874772.

Salama:1989:SDA

[131] Annika Salama and Mikko Nikinmaa. Species differences in the adrenergic responses of fish red cells: studies on whitefish, pikeperch, trout and carp. Fish Physiology and Biochemistry, 6(3):167–173, May 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF01874773.

Hickie:1989:IDC

[132] B. E. Hickie, D. G. Dixon, and J. F. Leatherland. The influence of the dietary carbohydrate: lipid ratio on the chronic toxicity of sodium pentachlorophenate to rainbow trout (Salmo gairdneri Richardson). Fish Physiology and Biochemistry, 6(3):175–185, May 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/BF01874774.

Fitch:1989:LDI

[133] Neil A. Fitch. Lactate dehydrogenase isozymes in the trunk and cardiac muscles of an Antarctic teleost fish, Notothenia neglecta Nybelin. Fish Physiology and Biochemistry, 6(3):187–195, May 1989. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF01874775.

Anonymous:1989:RAS

[134] Anonymous. Referees of articles submitted to the journal in 1988. Fish Physiology and Biochemistry, 6(3):197–198, May 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/BF01874776.

F:1989:TCR

[135] Leatherland J. F. and Hilton J. W. Thyroidal compensation in rainbow trout (Salmo gairdneri) fed canola meal. Fish Physiology and Biochemistry, 6(3):199, May 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01874777.

Celander:1989:CPM

[136] Malin Celander, Lars Förlin, and Tommy Andersson. Cytochrome P-450 mediated O-dealkylation of 7-alkoxycoumarins in liver microsomes from rainbow trout (Salmo gairdneri). Fish Physiology and Biochemistry, 6(4):199–205, July 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875022.

Asselt:1989:CID

[137] L. A. C. Van Asselt, R. J. M. Timmers, and P. G. W. J. Van Oordt. Catecholestrogens inhibit dopamine methylation in the gonadotrops of the African catfish, Clarias gariepinus. Fish Physiology and Biochemistry, 6(4):207–212, July 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875023.

Uematsu:1989:AIO

[138] Kazumasa Uematsu, Susanne Holmgren, and Stefan Nilsson. Autonomic innervation of the ovary of the Atlantic cod, Gadus morhua. Fish Physiology and Biochemistry, 6(4):213–219, July 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF01875024.

Olsson:1989:IMS

[139] Per-Erik Olsson, Åke Larsson, and Lashitew Gedamu. Induction of metallothionein synthesis in rainbow trout, Salmo gairdneri, during long-term exposure to waterborne cadmium. Fish Physiology and Biochemistry, 6(4):221–229, July 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875025.

McCormick:1989:MEK

[140] Stephen D. McCormick, Richard L. Saunders, and Arthur D. MacIntyre. Mitochondrial enzyme and Na^{+}, K^{+}-ATPase activity, and ion regulation during parr-smolt transformation of Atlantic salmon (Salmon salar). Fish Physiology and Biochemistry, 6(4):231–241, July 1989. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF01875026.

McCormick:1989:ISE

[141] Stephen D. McCormick, Christopher D. Moyes, and James S. Ballantyne. Influence of salinity on the energetics of gill and kidney of Atlantic salmon (Salmo salar). Fish Physiology and Biochemistry, 6(4):243–254, July 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF01875027.

Cyr:1989:EEV

[142] Daniel G. Cyr and J. G. Eales. T_{3} enhancement of in vitro estradiol-17β secretion by oocytes of rainbow trout, Salmo gairdneri, is dependent on cAMP. Fish Physiology and Biochemistry, 6(4):255–259, July 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF01875028.

Greenblatt:1989:CTH

[143] Mike Greenblatt, Christopher L. Brown, and Howard A. Bern. Changes in thyroid hormone levels in eggs and larvae and in iodide uptake by eggs of Coho and Chinook salmon, Oncorhynchus kisutch and O. tshawytscha. Fish Physiology and Biochemistry, 6(5):261–278, September 1989. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF01881680.

Kah:1989:GSB

[144] O. Kah, S. Dufour, and B. Breton. The GnRH systems in the brain and pituitary of normal and hCG treated European silver eels. Fish Physiology and Biochemistry, 6(5):279–284, September 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF01881681.

Lin:1989:DIG

[145] Hao-Ren Lin, Chun Peng, and Richard E. Peter. Dopamine inhibits gonadotropin secretion in the Chinese loach (Paramisgurnus dabryanus). Fish Physiology and Biochemistry, 6(5):285–288, September 1989. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF01881682.

Haider:1989:REV

[146] S. Haider and R. Moses Inbaraj. Relative in vitro effectiveness of estradiol-17β, androgens, corticosteroids, progesterone and other pregnene derivatives on germinal vesicle breakdown in oocytes of Indian major carps, Labeo rohita, Cirrhinus mrigala and Catla catla. Fish Physiology and Biochemistry, 6(5):289–295, September 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF01881683.

Wright:1989:CAG

[147] Patricia A. Wright and Steve F. Perry. A critical analysis of gas transfer in vivo and in a blood-perfused trout head preparation. Fish Physiology and Biochemistry, 6(5):297–308, September 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF01881684.

Altringham:1989:IPF

[148] J. D. Altringham and I. A. Johnston. The innervation pattern of fast myotomal muscle in the teleost Myoxocephalus scorpius: A reappraisal. Fish Physiology and Biochemistry, 6(5):309–313, September 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF01881685.

Axelsson:1989:RCO

[149] Michael Axelsson, William R. Driedzic, and Stefan Nilsson. Regulation of cardiac output and gut blood flow in the sea raven, Hemitripterus americanus. Fish Physiology and Biochemistry, 6(5):315–326, September 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF01881686.

Forster:1989:PHH

[150] M. E. Forster. Performance of the heat of the hagfish, Eptatretus cirrhatus. Fish Physiology and Biochemistry, 6(5):327–331, September 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF01881687.

vonderDecken:1989:MHC

[151] Alexandra von der Decken and Einar Lied. Myosin heavy chain synthesis in white trunk muscle of cod (Gadus morhua) fed different ration sizes. Fish Physiology and Biochemistry, 6(6):333–340, November 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF01875603.

Gordon:1989:ESF

[152] Malcolm S. Gordon, Helen G. Chin, and Marija Vojkovich. Energetics of swimming in fishes using different methods of locomotion: I. Labriform swimmers. Fish Physiology and Biochemistry, 6(6):341–352, November 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF01875604.

Johnston:1989:RBD

[153] W. L. Johnston, E. MacDonald, and J. W. Hilton. Relationships between dietary ascorbic acid status and deficiency, weight gain and brain neurotransmitter levels in juvenile rainbow trout, Salmo gairdneri. Fish Physiology and Biochemistry, 6(6):353–365, November 1989. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF01875605.

Fritzvold:1989:LPC

[154] Randi Fritzvold, Birgit H. Dannevig, and Trond Berg. The liver parenchymal cells of rainbow trout (Salmo gairdneri) endocytose mannoseterminated glycoproteins. Fish Physiology and Biochemistry, 6(6):367– 375, November 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875606.

Peter:1989:OMT

[155] M. C. Subash Peter and Oommen V. Oommen. Oxidative metabolism in a teleost, Anabas testudineus Bloch: effect of testosterone and estradiol-17 β on hepatic enzyme activities. Fish Physiology and Biochemistry, 6 (6):377–385, November 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875607.

Ottolenghi:1989:GAG

[156] C. Ottolenghi, A. C. Puviani, and E. M. Plisetskaya. Glycogenolytic action of glucagon-family peptides and epinephrine on catfish hepatocytes. Fish Physiology and Biochemistry, 6(6):387–394, November 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF01875608.

Shephard:1989:AIE

[157] K. L. Shephard. An analysis of the ion-exchange characteristics of fishegg chorions. Fish Physiology and Biochemistry, 6(6):395–401, November 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF01875609.

Hew:1989:TFP

[158] Choy L. Hew. Transgenic fish: present status and future directions. Fish Physiology and Biochemistry, 7(1-6):??, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/BF00004736.

Chen:1989:EIT

[159] T. T. Chen, L. B. Agellon, and D. A. Powers. Evolutionary implications of two rainbow trout growth hormone genes. Fish Physiology and Biochemistry, 7(1-6):??, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004732.

Marshall:1989:CIT

[160] William S. Marshall, Sharon E. Bryson, and David R. Idler. Control of ion transport by the sperm duct epithelium of brook trout (Salvelinus fontinalis). Fish Physiology and Biochemistry, 7(1-6):??, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004725.

Thomas:1989:ETP

[161] Peter Thomas and John M. Trant. Evidence that 17α, 20β, 21-trihydroxy-4-pregnen-3-one is a maturation-inducing steroid in spotted seatrout. Fish Physiology and Biochemistry, 7(1-6):??, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/BF00004706.

Plisetskaya:1989:PFE

[162] Erika M. Plisetskaya. Physiology of fish endocrine pancreas. Fish Physiology and Biochemistry, 7(1-6):??, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004688.

Pang:1989:P

[163] P. K. T. Pang, R. E. Peter, and M. P. Schreibman. Preface. Fish Physiology and Biochemistry, 7(1-6):1–2, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004683.

Gorbman:1989:RFE

[164] Aubrey Gorbman. Roots of fish endocrinology, a perspective. Fish Physiology and Biochemistry, 7(1-6):3–10, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004684.

Grau:1989:TPC

[165] E. Gordon Grau and Lisa M. H. Helms. The tilapia prolactin cell: A model for stimulus-secretion coupling. Fish Physiology and Biochemistry, 7(1-6):11–19, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004685.

Fryer:1989:NRA

[166] James N. Fryer. Neuropeptides regulating the activity of goldfish corticotropes and melanotropes. Fish Physiology and Biochemistry, 7 (1-6):21–27, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004686.

Kawauchi:1989:DTG

[167] Hiroshi Kawauchi, Kunimasa Suzuki, and Seiga Itoh. The duality of teleost gonadotropins. Fish Physiology and Biochemistry, 7(1-6):29–38, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004687.

Billard:1989:EFC

[168] R. Billard. Endocrinology and fish culture. Fish Physiology and Biochemistry, 7(1-6):49–58, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004689.

Zohar:1989:BGR

[169] Y. Zohar, A. Goren, and Y. Koch. The bioactivity of gonadotropin releasing hormones and its regulation in the gilthead seabream, Sparus aurata: in vivo and in vitro studies. Fish Physiology and Biochemistry, 7(1-6):59–67, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004690.

Kah:1989:CCD

[170] O. Kah, A. Pontet, and A. Calas. Characterization, cerebral distribution and gonadotropin release activity of neuropeptide Y (NPY) in the goldfish. Fish Physiology and Biochemistry, 7(1-6):69–76, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004691.

Breton:1989:NNM

[171] Bernard Breton, Tomasz Mikolajczyk, and Hubert Vaudry. Neuropeptide Y (NPY) modulates in vitro gonadotropin in release from rainbow trout pituitary glands. Fish Physiology and Biochemistry, 7(1-6):77–83, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004692.

Sherwood:1989:OMF

[172] Nancy M. Sherwood and David A. Lovejoy. The origin of the mammalian form of GnRH in primitive fishes. Fish Physiology and Biochemistry, 7(1-6):85–93, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004693.

Crim:1989:MST

[173] Laurence W. Crim, Donna M. Evans, and Kay Moreland. Monoaminergic substances in the teleost brain: Catecholamine levels in male and female winter flounder, Pseudopleuronectes americanus Walbaum, associated with gonadal recrudescence. Fish Physiology and Biochemistry, 7(1-6):95–100, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004694.

Sower:1989:ELG

[174] Stacia A. Sower. Effects of lamprey gonadotropin-releasing hormone and analogs on steroidogenesis and spermiation in male sea lampreys. Fish Physiology and Biochemistry, 7(1-6):101–107, June 1989. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004695.

Habibi:1989:PGR

[175] H. R. Habibi, R. De Leeuw, and R. E. Peter. Pituitary gonadotropinreleasing hormone (GnRH) receptor activity in goldfish and catfish: seasonal and gonadal effects. Fish Physiology and Biochemistry, 7(1-6):109–118, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004696.

Specker:1989:UAT

[176] Jennifer L. Specker, Patricia Stockin Brown, and Stephen C. Brown. Unequal activities of the two tilapia prolactins in the whole-animal transepithelial potential bioassay using the red eft. Fish Physiology and Biochemistry, 7(1-6):119–124, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004697.

Youson:1989:MDE

[177] John H. Youson and W. Mar Elliott. Morphogenesis and distribution of the endocrine pancreas in adult lampreys. Fish Physiology and Biochemistry, 7(1-6):125–131, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004698.

Marchant:1989:HPI

[178] Tracy A. Marchant and Richard E. Peter. Hypothalamic peptides influencing growth hormone secretion in the goldfish, Carassius auratus. Fish Physiology and Biochemistry, 7(1-6):133–139, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004699.

Kobayashi:1989:ISH

[179] Makito Kobayashi, Katsumi Aida, and Isao Hanyu. Involvement of steroid hormones in the preovulatory gonadotropin surge in female goldfish. Fish Physiology and Biochemistry, 7(1-6):141–146, June 1989. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004700.

Dickhoff:1989:RBM

[180] Walton W. Dickhoff, Liguang Yan, and Melinda G. Bernard. Relationship between metabolic and reproductive hormones in salmonid fish. Fish Physiology and Biochemistry, 7(1-6):147–155, June 1989. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004701.

Dufour:1989:PFC

[181] Sylvie Dufour, Nadine Le Belle, and Yves-Alain Fontaine. Positive feedback control by the gonads on gonadotropin (GTH) and gonadoliberin (GnRH) levels in experimentally matured female silver eels, Anguilla anguilla. Fish Physiology and Biochemistry, 7(1-6):157–162, June 1989. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004702.

Goetz:1989:RPC

[182] Frederick William Goetz, Priscilla Duman, and Erik G. Janowsky. The role of prostaglandins in the control of ovulation in yellow perch, Perca flavescens. Fish Physiology and Biochemistry, 7(1-6):163–168, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004703.

Swanson:1989:GIJ

[183] Penny Swanson, M. Bernard, and W. W. Dickhoff. Gonadotropins I and II in juvenile Coho salmon. Fish Physiology and Biochemistry, 7 (1-6):169–176, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004704.

Yu:1989:IPG

[184] J. Y. L. Yu and S. T. Shen. Isolation of pituitary glycoprotein gonadotropins from the grass carp (Ctenopharyngodon idell). Fish Physiology and Biochemistry, 7(1-6):177–183, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004705.

Nagahama:1989:MSA

[185] Yoshitaka Nagahama and Masakane Yamashita. Mechanisms of synthesis and action of 17α, 20β-dihydroxy-4-pregnen-3-one, a teleost maturation-inducing substance. Fish Physiology and Biochemistry, 7 (1-6):193–200, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004707.

Timmers:1989:CMB

[186] R. J. M. Timmers and J. G. D. Lambert. Catechol-o-methyltransferase in the brain of the male African catfish, Clarias gariepinus; distribution and significance for the metabolism of catecholestrogens and dopamine. Fish Physiology and Biochemistry, 7(1-6):201–210, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004708.

Nakamura:1989:DDL

[187] Masaru Nakamura and Yoshitaka Nagahama. Differentiation and development of Leydig cells, and changes of testosterone levels during testicular differentiation in tilapia Oreochromis niloticus. Fish Physiology and Biochemistry, 7(1-6):211–219, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004709.

Dubois:1989:RSC

[188] Wilfrid Dubois and Gloria V. Callard. Role of the Sertoli cell in spermatogenesis: The Squalus testis model. Fish Physiology and Biochemistry, 7 (1-6):221–227, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004710.

Chan:1989:SSI

[189] S. T. H. Chan and W. S. B. Yeung. Sex steroids in intersexual fishes. Fish Physiology and Biochemistry, 7(1-6):229–235, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004711.

Schreibman:1989:GIR

[190] Martin P. Schreibman, Seymour Holtzman, and Ronald A. Eckhardt. Genetic influences on reproductive system development and function: A review. Fish Physiology and Biochemistry, 7(1-6):237–242, June 1989. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004712.

Bail:1989:PCS

[191] Pierre-Yves Le Bail, Geneviève Boulard, and Michel Zygmunt. Purification of Chinook salmon (Oncorhynchus tshawytscha) GH for receptor study. Fish Physiology and Biochemistry, 7(1-6):243–251, June 1989. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004713.

Pickering:1989:SRD

[192] A. D. Pickering and T. G. Pottinger. Stress responses and disease resistance in salmonid fish: Effects of chronic elevation of plasma cortisol. Fish Physiology and Biochemistry, 7(1-6):253–258, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004714.

Schreck:1989:RIF

[193] Carl B. Schreck, C. Samue Bradford, and Reynaldo Patiño. Regulation of the interrenal of fishes: non-classical control mechanisms. Fish Physiology and Biochemistry, 7(1-6):259–265, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004715.

Weisbart:1989:EAE

[194] M. Weisbart, P. K. Chakraborti, and H. C. Freeman. Effect of acid environments on cortisol and cortisol receptor activity in Atlantic salmon, Salmo salar. Fish Physiology and Biochemistry, 7(1-6):267–272, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004716.

Epple:1989:CCE

[195] August Epple, Charles B. Hathaway, and Barbara Nibbio. Circulatory catecholamines in the eel: origins and functions. Fish Physiology and Biochemistry, 7(1-6):273–278, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004717.

Mommsen:1989:MAG

[196] Thomas P. Mommsen and Thomas W. Moon. Metabolic actions of glucagon-family hormones in liver. Fish Physiology and Biochemistry, 7(1-6):279–288, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004718.

Eales:1989:RBP

[197] J. G. Eales and Deborah L. MacLatchy. The relationship between T_{3} production and energy balance in salmonids and other teleosts. Fish Physiology and Biochemistry, 7(1-6):289–293, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004719.

Brown:1989:ESS

[198] Christopher L. Brown, Serge I. Doroshov, and Howard A. Bern. Enhanced survival in striped bass fingerlings after maternal triiodothyronine treatment. Fish Physiology and Biochemistry, 7(1-6):295–299, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004720.

Gonnet:1989:HCP

[199] Francis Gonnet, Alain Barret, and Patrick Prunet. Hypothalamic control of prolactin release in the rainbow trout, Salmo gairdneri: in vitro studies. Fish Physiology and Biochemistry, 7(1-6):301–308, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004721.

Ogasawara:1989:CPP

[200] Tsuyoshi Ogasawara, Tetsuya Hirano, and Masatomo Tagawa. Changes in plasma prolactin and growth hormone concentrations during freshwater adaptation of juvenile chum salmon (Oncorhynchus keta) reared in seawater for a prolonged period. Fish Physiology and Biochemistry, 7 (1-6):309–313, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004722.

Collie:1989:SSS

[201] Nathan L. Collie, Jonathan P. Bolton, and Tetsuya Hirano. Survival of salmonids in seawater and the time-frame of growth hormone action. Fish Physiology and Biochemistry, 7(1-6):315–321, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004723.

Nishimura:1989:SCR

[202] Hiroko Nishimura and Margaret A. Madey. Signals controlling renin release in aglomerular toadfish. Fish Physiology and Biochemistry, 7 (1-6):323–329, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004724.

Kaneko:1989:LCR

[203] Toyoji Kaneko, Stephen Harvey, and Peter K. T. Pang. Localization of calcium regulatory hormones in fish. Fish Physiology and Biochemistry, 7(1-6):337–342, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004726.

Flik:1989:STC

[204] Gert Flik, Teresa Labedz, and Peter K. T. Pang. Studies on teleost corpuscles of Stannius: Physiological and biochemical aspects of synthesis and release of hypocalcin in trout, goldfish and eel. Fish Physiology and Biochemistry, 7(1-6):343–349, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004727.

Fenwick:1989:VDR

[205] J. C. Fenwick and M. G. Vermette. Vitamin D_{3} and the renal handling of phosphate in American eels. Fish Physiology and Biochemistry, 7(1-6):351–358, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004728.

Butkus:1989:PBC

[206] Aldona Butkus, Neale A. Yates, and John P. Coghlan. Processing and bioactivity of the corpuscles of Stannius protein of the Australian eel. Fish Physiology and Biochemistry, 7(1-6):359–365, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004729.

Wagner:1989:SSP

[207] Graham F. Wagner and Henry G. Friesen. Studies on the structure and physiology of salmon teleocalcin. Fish Physiology and Biochemistry, 7 (1-6):367–374, June 1989. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004730.